# Simplifying, Stabilizing & Scaling Continuous-Time Consistency Models

### v1

@@ 2.1 Diffusion Models
Given a training dataset, let $p_{d}$ denote its underlying data distribution and $\sigma_{d}$ its standard deviation. Diffusion models generate samples by learning to reverse a noising process that progressively perturbs a data sample ${\bm{x}}_{0}\sim p_{d}$ into a noisy version ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{z}}_{t}$, where ${\bm{z}}_{t}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$ is standard Gaussian noise. This perturbation increases with $t\in[0,T]$, where larger $t$ indicates greater noise.
@@
EDM (Karras et al., 2022; 2024).   The noising process simply sets $\alpha_{t}=1$ and $\sigma_{t}=t$, with the training objective given by $\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\left\|{\bm{f}}_{\theta}({\bm{x}}_{t},t)-{\bm{x}}_{0}\right\|_{2}^{2}\right]$, where $w(t)$ is a weighting function. The diffusion model is parameterized as ${\bm{f}}_{\theta}({\bm{x}}_{t},t)=c_{\text{skip}}(t){\bm{x}}_{t}+c_{\text{out}}(t){\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$, and $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$, and $c_{\text{noise}}$ are manually designed coefficients that ensure the training objective has the unit variance across timesteps at initialization. For sampling, EDM solves the probability flow ODE (PF-ODE)  (Song et al., 2021b), defined by $\frac{\mathrm{d}{\bm{x}}_{t}}{\mathrm{d}t}=[{\bm{x}}_{t}-{\bm{f}}_{\theta}({\bm{x}}_{t},t)]/t$, starting from ${\bm{x}}_{T}\sim{\mathcal{N}}({\bm{0}},T^{2}{\bm{I}})$ and stopping at ${\bm{x}}_{0}$.
@@ 2.2 Consistency Models
Continuous-time CMs. When using $d({\bm{x}},{\bm{y}})=\|{\bm{x}}-{\bm{y}}\|_{2}^{2}$ and taking the limit $\Delta t\to 0$, Song et al. (2023) show that the gradient of Eq. (1) with respect to $\theta$ converges to
@@ 3 Simplifying Continuous-Time Consistency Models
Previous consistency models (CMs) adopt the model parameterization and diffusion process formulation in EDM (Karras et al., 2022). Specifically, the CM is parameterized as ${\bm{f}}_{\theta}({\bm{x}}_{t},t)=c_{\text{skip}}(t){\bm{x}}_{t}+c_{\text{out}}(t){\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$. The coefficients $c_{\text{skip}}(t)$, $c_{\text{out}}(t)$, $c_{\text{in}}(t)$ are fixed to ensure that the variance of the diffusion objective is equalized across all time steps at initialization, and $c_{\text{noise}}(t)$ is a transformation of $t$ for better time conditioning. Since EDM diffusion process is variance-exploding (Song et al., 2021b), meaning that ${\bm{x}}_{t}={\bm{x}}_{0}+t{\bm{z}}_{t}$, we can derive that $c_{\text{skip}}(t)=\sigma_{d}^{2}/(t^{2}+\sigma_{d}^{2})$, $c_{\text{out}}(t)=\sigma_{d}\cdot t/\sqrt{\sigma_{d}^{2}+t^{2}}$, and $c_{\text{in}}(t)=1/\sqrt{t^{2}+\sigma_{d}^{2}}$ (see Appendix B.6 in Karras et al. (2022)). Although these coefficients are important for training efficiency, their complex arithmetic relationships with $t$ and $\sigma_{d}$ complicate theoretical analyses of CMs.
To simplify EDM and subsequently CMs, we propose TrigFlow, a formulation of diffusion models that keep the EDM properties but satisfy $c_{\text{skip}}(t)=\cos(t)$, $c_{\text{out}}(t)=\sin(t)$, and $c_{\text{in}}(t)\equiv 1/\sigma_{d}$ (proof in Appendix B). TrigFlow is a special case of flow matching (also known as stochastic interpolants or rectified flows) and v-prediction parameterization (Salimans & Ho, 2022). It closely resembles the trigonometric interpolant proposed by Albergo & Vanden-Eijnden (2023); Albergo et al. (2023); Ma et al. (2024), but is modified to account for $\sigma_{d}$, the standard deviation of the data distribution $p_{d}$. Since TrigFlow is a special case of flow matching and simultaneously satisfies EDM principles, it combines the advantages of both formulations while allowing the diffusion process, diffusion model parameterization, the PF-ODE, the diffusion training objective, and the CM parameterization to all have simple expressions, as provided below.
@@
Diffusion Models and PF-ODE. We parameterize the diffusion model as ${\bm{f}}_{\theta}({\bm{x}}_{t},t)={\bm{F}}_{\theta}({\bm{x}}_{t}/\sigma_{d},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$, and $c_{\text{noise}}(t)$ is a transformation of $t$ to facilitate time conditioning. The corresponding PF-ODE is given by
@@ 4.1 Parameterization and Network Architecture
Identity Time Transformation ($\bm{c_{\text{noise}}(t)=t}$). Most existing CMs use the EDM formulation, which can be directly translated to the TrigFlow formulation as described in Appendix B.2. In particular, the time transformation becomes $c_{\text{noise}}(t)=\log(\sigma_{d}\tan t)$. Straightforward derivation shows that with this $c_{\text{noise}}(t)$, $\sin(t)\cdot\partial_{t}c_{\text{noise}}(t)=1/\cos(t)$ blows up whenever $t\to\frac{\pi}{2}$. To mitigate numerical instability, we propose to use $c_{\text{noise}}(t)=t$ as the default time transformation.
@@
Adaptive Double Normalization. Song & Dhariwal (2023) found that the AdaGN layer (Dhariwal & Nichol, 2021), defined as ${\bm{y}}=\text{norm}({\bm{x}})\odot{\bm{s}}(t)+{\bm{b}}(t)$, negatively impacts CM training. Our modification is adaptive double normalization, defined as ${\bm{y}}=\text{norm}({\bm{x}})\odot\text{pnorm}({\bm{s}}(t))+\text{pnorm}({\bm{b}}(t))$, where $\text{pnorm}(\cdot)$ denotes pixel normalization (Karras, 2017). Empirically we find it retains the expressive power of AdaGN for diffusion training but removes its instability in CM training.
@@ 4.2 Training Objectives
Diffusion Finetuning and Tangent Warmup. For consistency distillation, we find that finetuning the CM from a pretrained diffusion model can speed up convergence, which is consistent with Song et al. (2023); Geng et al. (2024). Recall that in Eq. (6), the tangent $\frac{\mathrm{d}{\bm{f}}_{\theta^{-}}}{\mathrm{d}t}$ can be decomposed into two parts: the first term $\cos(t)(\sigma_{d}{\bm{F}}_{\theta^{-}}-\frac{\mathrm{d}{\bm{x}}_{t}}{\mathrm{d}t})$ is relatively stable, whereas the second term $\sin(t)({\bm{x}}_{t}+\sigma_{d}\frac{\mathrm{d}{\bm{F}}_{\theta^{-}}}{\mathrm{d}t})$ causes instability. To alleviate this problem, we gradually warm up the second term by replacing the coefficient $\sin(t)$ with $r\cdot\sin(t)$, where $r$ linearly increases from 00 to $1$ over the first 10k training iterations.
@@ 6 Conclusion
Acknowledgements
We would like to thank Allan Jabri, Aaron Lou, Alex Nichol, Huiwen Chang, Heewoo Jun, and Ishaan Gulrajani for technical discussions, and Mingyu Ye for assistance with plots and diagrams. We also appreciate the support from Mark Chen and Prafulla Dhariwal for this research project.
@@ B.1 Derivations
${\mathcal{L}}_{\text{Diff}}=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)c_{\text{out}}^{2}(t)\left\|{\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))-\frac{c_{\text{skip}}(t)\alpha_{t}{\bm{x}}_{0}-(1-c_{\text{skip}}(t)\sigma_{t}){\bm{z}}}{c_{\text{out}}(t)}\right\|_{2}^{2}\right].$ (11)
@@
$c_{\text{out}}^{2}(t)=\sigma_{d}^{2}c_{\text{skip}}^{2}(t)\alpha_{t}^{2}+(c_{\text{skip}}(t)\sigma_{t}-1)^{2}.$ (13)
@@
$\begin{split}c_{\text{skip}}(t)=\frac{\sigma_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}},\quad c_{\text{out}}(t)=\pm\frac{\sigma_{d}\alpha_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}.\end{split}$ (14)
Though equivalent, we choose $c_{\text{out}}(t)=-\frac{\sigma_{d}\alpha_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}$ which can simplify some derivations below.
@@
${\bm{D}}_{\theta}({\bm{x}}_{t},t)$ $=\frac{\sigma_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}}{\bm{x}}_{t}-\frac{\alpha_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}\sigma_{d}{\bm{F}}_{\theta}\left(\frac{{\bm{x}}_{t}}{\sigma_{d}\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},c_{\text{noise}}(t)\right),$ (15)
${\mathcal{L}}_{\text{Diff}}$ $=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\frac{\alpha^{2}_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}}\left\|\sigma_{d}{\bm{F}}_{\theta}\left(\frac{{\bm{x}}_{t}}{\sigma_{d}\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},c_{\text{noise}}(t)\right)-\frac{\alpha_{t}{\bm{z}}-\sigma_{t}{\bm{x}}_{0}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}\right\|_{2}^{2}\right].$ (16)
@@
${\bm{D}}_{\theta}({\bm{x}}_{t},t)$ $=\hat{\sigma}_{t}\hat{{\bm{x}}}_{t}-\hat{\alpha}_{t}\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{t}}{\sigma_{d}},c_{\text{noise}}(t)\right),$ (17)
${\mathcal{L}}_{\text{Diff}}$ $=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\hat{\alpha}_{t}^{2}\left\|\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{t}}{\sigma_{d}},c_{\text{noise}}(t)\right)-(\hat{\alpha}_{t}{\bm{z}}-\hat{\sigma}_{t}{\bm{x}}_{0})\right\|_{2}^{2}\right].$ (18)
@@
$\hat{{\bm{x}}}_{t}=\frac{\hat{\sigma}_{t}}{\hat{\sigma}_{s}}\hat{{\bm{x}}}_{s}+\hat{\sigma}_{t}\int_{\hat{\lambda}_{s}}^{\hat{\lambda}_{t}}e^{\hat{\lambda}}{\bm{D}}_{\theta}(\hat{{\bm{x}}}_{\hat{\lambda}},\hat{\lambda})\mathrm{d}\hat{\lambda},$ (20)
which is exactly the sampling procedure of the diffusion process $\hat{{\bm{x}}_{t}}$, which means noise schedules of diffusion models won’t affect the performance of sampling. In other words, for any diffusion process $(\alpha_{t},\sigma_{t},{\bm{x}}_{t})$ at time $t$, we can always divide them by $\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}$ to obtain the diffusion process $(\hat{\alpha}_{t},\hat{\sigma}_{t},\hat{{\bm{x}}}_{t})$ with $\hat{\alpha}_{t}^{2}+\hat{\sigma}_{t}^{2}=1$ and all the parameterization, training objective and sampling procedure can be equivalently transformed. The only difference is the corresponding training weighting $w(t)\sigma^{2}_{d}\hat{\alpha}_{t}^{2}$ in Eq. (18), which we will discuss in the next step.
A straightforward corollary is that the “optimal transport path” (Lipman et al., 2022) in flow matching with $\alpha_{t}=1-t,\sigma_{t}=t$ can be equivalently converted to other noise schedules. The reason of its better empirical performance is essentially due to the different weighting during training and the lack of advanced diffusion sampler such as DPM-Solver series (Lu et al., 2022a; b) during sampling, not the “straight path” (Lipman et al., 2022) itself.
@@
${\mathcal{L}}_{\text{Diff}}=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}}}\left[\int_{0}^{\frac{\pi}{2}}\underbrace{p\left(\hat{t}\right)w\left(\hat{t}\right)\cos^{2}\left(\hat{t}\right)}_{\text{training weighting}}\left\|\underbrace{\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{\hat{t}}}{\sigma_{d}},c_{\text{noise}}\left(\hat{t}\right)\right)-(\cos\left(\hat{t}\right){\bm{z}}-\sin\left(\hat{t}\right){\bm{x}}_{0})}_{\text{independent from $\alpha_{t}$ and $\sigma_{t}$}}\right\|_{2}^{2}\mathrm{d}\hat{t}\right].$ (22)

### v2

@@ 2.1 Diffusion Models
Given a training dataset, let $p_{d}$ denote its underlying data distribution and $\sigma_{d}$ its standard deviation. Diffusion models generate samples by learning to reverse a noising process that progressively perturbs a data sample ${\bm{x}}_{0}\sim p_{d}$ into a noisy version ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{z}}$, where ${\bm{z}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$ is standard Gaussian noise. This perturbation increases with $t\in[0,T]$, where larger $t$ indicates greater noise.
@@
EDM (Karras et al., 2022; 2024).   The noising process simply sets $\alpha_{t}=1$ and $\sigma_{t}=t$, with the training objective given by $\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\left\|{\bm{f}}^{\text{DM}}_{\theta}({\bm{x}}_{t},t)-{\bm{x}}_{0}\right\|_{2}^{2}\right]$, where $w(t)$ is a weighting function. The diffusion model is parameterized as ${\bm{f}}^{\text{DM}}_{\theta}({\bm{x}}_{t},t)=c_{\text{skip}}(t){\bm{x}}_{t}+c_{\text{out}}(t){\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$, and $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$, and $c_{\text{noise}}$ are manually designed coefficients that ensure the training objective has the unit variance across timesteps at initialization. For sampling, EDM solves the probability flow ODE (PF-ODE)  (Song et al., 2021b), defined by $\frac{\mathrm{d}{\bm{x}}_{t}}{\mathrm{d}t}=[{\bm{x}}_{t}-{\bm{f}}^{\text{DM}}_{\theta}({\bm{x}}_{t},t)]/t$, starting from ${\bm{x}}_{T}\sim{\mathcal{N}}({\bm{0}},T^{2}{\bm{I}})$ and stopping at ${\bm{x}}_{0}$.
@@ 2.2 Consistency Models
Continuous-time CMs. When using $d({\bm{x}},{\bm{y}})=\|{\bm{x}}-{\bm{y}}\|_{2}^{2}$ and taking the limit $\Delta t\to 0$, Song et al. (2023, Remark 10) show that the gradient of Eq. (1) with respect to $\theta$ converges to
@@ 3 Simplifying Continuous-Time Consistency Models
Previous consistency models (CMs) adopt the model parameterization and diffusion process formulation in EDM (Karras et al., 2022). Specifically, the CM is parameterized as ${\bm{f}}_{\theta}({\bm{x}}_{t},t)=c_{\text{skip}}(t){\bm{x}}_{t}+c_{\text{out}}(t){\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$. The coefficients $c_{\text{skip}}(t)$, $c_{\text{out}}(t)$, $c_{\text{in}}(t)$ are fixed to ensure that the variance of the diffusion objective is equalized across all time steps at initialization, and $c_{\text{noise}}(t)$ is a transformation of $t$ for better time conditioning. Since EDM diffusion process is variance-exploding (Song et al., 2021b), meaning that ${\bm{x}}_{t}={\bm{x}}_{0}+t{\bm{z}}$, we can derive that $c_{\text{skip}}(t)=\sigma_{d}^{2}/(t^{2}+\sigma_{d}^{2})$, $c_{\text{out}}(t)=\sigma_{d}\cdot t/\sqrt{\sigma_{d}^{2}+t^{2}}$, and $c_{\text{in}}(t)=1/\sqrt{t^{2}+\sigma_{d}^{2}}$ (see Appendix B.6 in Karras et al. (2022)). Although these coefficients are important for training efficiency, their complex arithmetic relationships with $t$ and $\sigma_{d}$ complicate theoretical analyses of CMs.
To simplify EDM and subsequently CMs, we propose TrigFlow, a formulation of diffusion models that keep the EDM properties but satisfy $c_{\text{skip}}(t)=\cos(t)$, $c_{\text{out}}(t)=-\sigma_{d}\sin(t)$, and $c_{\text{in}}(t)\equiv 1/\sigma_{d}$ (proof in Appendix B). TrigFlow is a special case of flow matching (also known as stochastic interpolants or rectified flows) and v-prediction parameterization (Salimans & Ho, 2022). It closely resembles the trigonometric interpolant proposed by Albergo & Vanden-Eijnden (2023); Albergo et al. (2023); Ma et al. (2024), but is modified to account for $\sigma_{d}$, the standard deviation of the data distribution $p_{d}$. Since TrigFlow is a special case of flow matching and simultaneously satisfies EDM principles, it combines the advantages of both formulations while allowing the diffusion process, diffusion model parameterization, the PF-ODE, the diffusion training objective, and the CM parameterization to all have simple expressions, as provided below.
@@
Diffusion Models and PF-ODE. We parameterize the diffusion model as ${\bm{f}}^{\text{DM}}_{\theta}({\bm{x}}_{t},t)={\bm{F}}_{\theta}({\bm{x}}_{t}/\sigma_{d},c_{\text{noise}}(t))$, where ${\bm{F}}_{\theta}$ is a neural network with parameters $\theta$, and $c_{\text{noise}}(t)$ is a transformation of $t$ to facilitate time conditioning. The corresponding PF-ODE is given by
@@ 4.1 Parameterization and Network Architecture
Identity Time Transformation ($\bm{c_{\text{noise}}(t)=t}$). Most existing CMs use the EDM formulation, which can be directly translated to the TrigFlow formulation as described in Appendix B.2. In particular, the time transformation becomes $c_{\text{noise}}(t)\propto\log(\sigma_{d}\tan t)$. Straightforward derivation shows that with this $c_{\text{noise}}(t)$, $\sin(t)\cdot\partial_{t}c_{\text{noise}}(t)=1/\cos(t)$ blows up whenever $t\to\frac{\pi}{2}$. To mitigate numerical instability, we propose to use $c_{\text{noise}}(t)=t$ as the default time transformation.
@@
Adaptive Double Normalization. Song & Dhariwal (2023) found that the AdaGN layer (Dhariwal & Nichol, 2021), defined as ${\bm{y}}=\text{norm}({\bm{x}})\odot{\bm{s}}(t)+{\bm{b}}(t)$, negatively causes CM training to diverge. Our modification is adaptive double normalization, defined as ${\bm{y}}=\text{norm}({\bm{x}})\odot\text{pnorm}({\bm{s}}(t))+\text{pnorm}({\bm{b}}(t))$, where $\text{pnorm}(\cdot)$ denotes pixel normalization (Karras, 2017). Empirically we find it retains the expressive power of AdaGN for diffusion training but removes its instability in CM training.
@@ 4.2 Training Objectives
Diffusion Finetuning and Tangent Warmup. For consistency distillation, we find that finetuning the CM from a pretrained diffusion model can speed up convergence, which is consistent with Song et al. (2023); Geng et al. (2024). Recall that in Eq. (6), the tangent $\frac{\mathrm{d}{\bm{f}}_{\theta^{-}}}{\mathrm{d}t}$ can be decomposed into two parts: the first term $\cos(t)(\sigma_{d}{\bm{F}}_{\theta^{-}}-\frac{\mathrm{d}{\bm{x}}_{t}}{\mathrm{d}t})$ is relatively stable, whereas the second term $\sin(t)({\bm{x}}_{t}+\sigma_{d}\frac{\mathrm{d}{\bm{F}}_{\theta^{-}}}{\mathrm{d}t})$ may cause instability. We introduce an optional technique named as tangent warmup by replacing the coefficient $\sin(t)$ with $r\cdot\sin(t)$, where $r$ linearly increases from 00 to $1$ over the first 10k training iterations. We find that the tangent normalization does not affect sample quality but may reduce some gradient spikes during training.
@@ B.1 Derivations
${\mathcal{L}}_{\text{Diff}}=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)c_{\text{out}}^{2}(t)\left\|{\bm{F}}_{\theta}(c_{\text{in}}(t){\bm{x}}_{t},c_{\text{noise}}(t))-\frac{(1-c_{\text{skip}}(t)\alpha_{t}){\bm{x}}_{0}-c_{\text{skip}}(t)\sigma_{t}{\bm{z}}}{c_{\text{out}}(t)}\right\|_{2}^{2}\right].$ (11)
@@
$c_{\text{out}}^{2}(t)=\sigma_{d}^{2}(1-c_{\text{skip}}(t)\alpha_{t})^{2}+\sigma_{d}^{2}c_{\text{skip}}^{2}(t)\sigma_{t}^{2}.$ (13)
@@
$\begin{split}c_{\text{skip}}(t)=\frac{\alpha_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}},\quad c_{\text{out}}(t)=\pm\frac{\sigma_{d}\sigma_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}.\end{split}$ (14)
Though equivalent, we choose $c_{\text{out}}(t)=-\frac{\sigma_{d}\sigma_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}$ which can simplify some derivations below.
@@
${\bm{D}}_{\theta}({\bm{x}}_{t},t)$ $=\frac{\alpha_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}}{\bm{x}}_{t}-\frac{\sigma_{t}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}\sigma_{d}{\bm{F}}_{\theta}\left(\frac{{\bm{x}}_{t}}{\sigma_{d}\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},c_{\text{noise}}(t)\right),$ (15)
${\mathcal{L}}_{\text{Diff}}$ $=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\frac{\sigma^{2}_{t}}{\alpha_{t}^{2}+\sigma_{t}^{2}}\left\|\sigma_{d}{\bm{F}}_{\theta}\left(\frac{{\bm{x}}_{t}}{\sigma_{d}\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}},c_{\text{noise}}(t)\right)-\frac{\alpha_{t}{\bm{z}}-\sigma_{t}{\bm{x}}_{0}}{\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}}\right\|_{2}^{2}\right].$ (16)
@@
${\bm{D}}_{\theta}({\bm{x}}_{t},t)$ $=\hat{\alpha}_{t}\hat{{\bm{x}}}_{t}-\hat{\sigma}_{t}\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{t}}{\sigma_{d}},c_{\text{noise}}(t)\right),$ (17)
${\mathcal{L}}_{\text{Diff}}$ $=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}},t}\left[w(t)\hat{\sigma}_{t}^{2}\left\|\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{t}}{\sigma_{d}},c_{\text{noise}}(t)\right)-(\hat{\alpha}_{t}{\bm{z}}-\hat{\sigma}_{t}{\bm{x}}_{0})\right\|_{2}^{2}\right].$ (18)
@@
$\hat{{\bm{x}}}_{t}=\frac{\hat{\sigma}_{t}}{\hat{\sigma}_{s}}\hat{{\bm{x}}}_{s}+\hat{\sigma}_{t}\int_{\hat{\lambda}_{s}}^{\hat{\lambda}_{t}}e^{\hat{\lambda}}\hat{{\bm{D}}}_{\theta}(\hat{{\bm{x}}}_{\hat{\lambda}},\hat{\lambda})\mathrm{d}\hat{\lambda},$ (20)
which is exactly the sampling procedure of the diffusion process $\hat{{\bm{x}}_{t}}$, which means noise schedules of diffusion models won’t affect the performance of sampling. In other words, for any diffusion process $(\alpha_{t},\sigma_{t},{\bm{x}}_{t})$ at time $t$, we can always divide them by $\sqrt{\alpha_{t}^{2}+\sigma_{t}^{2}}$ to obtain the diffusion process $(\hat{\alpha}_{t},\hat{\sigma}_{t},\hat{{\bm{x}}}_{t})$ with $\hat{\alpha}_{t}^{2}+\hat{\sigma}_{t}^{2}=1$ and all the parameterization, training objective and sampling procedure can be equivalently transformed. The only difference is the corresponding training weighting $w(t)\sigma^{2}_{d}\hat{\sigma}_{t}^{2}$ in Eq. (18), which we will discuss in the next step.
A straightforward corollary is that the “optimal transport path” (Lipman et al., 2022) in flow matching with $\alpha_{t}=1-t,\sigma_{t}=t$ can be equivalently converted to other noise schedules. The reason of its better empirical performance is essentially due to the different weighting during training and the lack of advanced diffusion sampler such as DPM-Solver series (Lu et al., 2022a; b) during sampling, not the “straight path” (Lipman et al., 2022) itself. By converting the diffusion process to the space satisfying $\sqrt{\hat{\alpha}_{t}^{2}+\hat{\sigma}_{t}^{2}=1}$, the path connection ${\bm{x}}_{0}$ and ${\bm{z}}$ has consistent variance which matches the unit-variance design principles of EDM.
@@
${\mathcal{L}}_{\text{Diff}}=\mathbb{E}_{{\bm{x}}_{0},{\bm{z}}}\left[\int_{0}^{\frac{\pi}{2}}\underbrace{p\left(\hat{t}\right)w\left(\hat{t}\right)\sin^{2}\left(\hat{t}\right)}_{\text{training weighting}}\left\|\underbrace{\sigma_{d}{\bm{F}}_{\theta}\left(\frac{\hat{{\bm{x}}}_{\hat{t}}}{\sigma_{d}},c_{\text{noise}}\left(\hat{t}\right)\right)-(\cos\left(\hat{t}\right){\bm{z}}-\sin\left(\hat{t}\right){\bm{x}}_{0})}_{\text{independent from $\alpha_{t}$ and $\sigma_{t}$}}\right\|_{2}^{2}\mathrm{d}\hat{t}\right].$ (22)